OpenStudy (anonymous):

What is the geometric pair of radical 8 and radical 2?

Directrix (directrix):

Does "geometric pair" mean "geometric mean?" The geometric mean would be 4 but I don't know if that's what you are seeking.

OpenStudy (anonymous):

Yes, geometric mean. Sorry. Haha. Thank you!

Directrix (directrix):

Okay. x is the geometric mean between two numbers means, in this case, that radical 8 is to x as x is to radical 2. √ 8 / x = x / √ 2 x^2 = √ 16 x = 4 Note: x also = -4 but by definition, geometric mean is not negative.

OpenStudy (anonymous):

Thanks! You can explain it better than my teacher. Haha.

Directrix (directrix):

Thanks for the compliment (and the medal).